# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v3

@@ Abstract
The foundation model enables general-purpose problem-solving and enjoys desirable rapid adaptation due to its adopted cross-task generalization paradigms, e.g., pretraining, meta-training, and finetuning. Recent advances in these paradigms show the crucial role of challenging tasks’ prioritized sampling in enhancing adaptation robustness. However, ranking task difficulties exhausts massive task queries to evaluate, thus computation and annotation intensive, which is typically unaffordable in practice. This work underscores the criticality of both adaptation robustness and learning efficiency, especially in scenarios where tasks are risky or costly to evaluate, e.g., policy evaluations in Markov decision processes (MDPs) or inference with large models. To this end, we present Model Predictive Task Sampling (MPTS) to establish connections between the task space and adaptation risk landscape to form a theoretical guideline in robust active task sampling. MPTS characterizes the task episodic information with a generative model and directly predicts task-specific adaptation risk values from posterior inference. The developed risk learner can amortize expensive evaluation and provably approximately rank task difficulties in the pursuit of task robust adaptation. MPTS can be seamlessly integrated into zero-shot, few-shot, and many-shot learning paradigms. Extensive experimental results are conducted to exhibit the superiority of the proposed framework, remarkably increasing task adaptation robustness and retaining learning efficiency in contrast to existing state-of-the-art (SOTA) methods. The code is available at the project site https://github.com/thu-rllab/MPTS.
@@ 1 Introduction
Figure 1: Framework of Model Predictive Task Sampling in Any-Shot Learning. a. The task episodic learning might involve MDPs, large models, or annotators in task construction for optimization. The adaptive machine learner can be a large model or RL policy. MPTS contains posterior inference, adaptation outcome simulation, and task subset selection for any-shot learning. [Snow: freeze models; Fire: update models] b. MPTS amortizes the process, e.g., task batch construction, environment interaction, or large models’ inference at a lower computational budget with the risk learner. c. The risk learner takes in the coupled identifier and the exact adaptation risk to train under streaming variational inference. d. The risk learner simulates the adaptation outcomes of $\hat{\mathcal{B}}$ candidate identifiers, computes acquisition scores, and selects $-\mathcal{B}$ ones for next iteration.
Enhancing machine learning models’ generalization across various scenarios is a long-standing pursuit in artificial general intelligence. The rise of generative artificial intelligence provides a promising solution to this problem, which facilitates the development of foundation models 1, 2, 3. Unlike traditional models, which are specific to learning tasks and cannot sufficiently meet the demands of time-sensitive responses, e.g., online intelligent recommendation or consultancy, the foundation model exhibits the advantages of fast deployment in a wide range of tasks without learning from scratch.
One secret behind foundation models’ success lies in rapid adaptation acquired from pretraining, meta-training, and finetuning. Take meta-learning as an example; it optimizes the machine learner over the task distribution and shares a similar cognition in human’s developing skills 4, where we consolidate cumulated experience into knowledge and quickly convert the few-shot guidelines to a specific skill. Technically, adaptation learning paradigms update machine learners episodically and leverage extracted problem-solving prior to unseen but related scenarios in a zero-shot, few-shot, or many-shot manner 1, 5. Such a mechanism exploits slow learning on massive tasks to solve new tasks fast at test time.
@@
Currently, there have been several studies in enhancing adaptation, e.g., the neural scaling law in foundation models 6, 7 implies that cross-task generalization scales with model complexity, task capacity, and computational power. Recognized as a few-shot learner, large language models (LLMs) mandate the collection of large datasets as tasks to train for generalization 8. In robotics, a popular schema for generalist policies is diversifying the decision-making environments, e.g., Markov decision processes (MDPs), for the robot to interact and execute policy optimization 9. As a consequence, task episodic learning is commonly used in learning to adapt over task distributions, where the task batch is randomly resampled, e.g., from a uniform random sampler, to optimize over iteration (see Fig. 1/2b). Meanwhile, these distributions are induced by the task identifiers; e.g., in Fig. 2a, diverse physics parameters are used to create diverse MDPs in domain randomization (DR) 9 and meta reinforcement learning (Meta-RL) 10.
Meanwhile, as distribution shifts 11 exist ubiquitously at the task level in the real world, adaptation robustness is catching increasing attention in the field 12, 13. Some trivial yet frequently adopted sampling strategies, e.g., random sampling from a uniform task distribution, are prone to encounter catastrophic failure in adaptation to risk-sensitive scenarios due to undersampling critical tasks. For example, traffic accidents are more informative than usual cases when training autonomous driving systems but occupy a small portion 14. This inspires some advances in challenging tasks’ prioritized sampling strategies 14, 15, 16, which necessitates task difficulty ranking during optimization. For the sake of improving adaptation robustness, these approaches 17, 18, 16, 19, 20 rank the task samples after evaluation and prioritize difficult ones to optimize (see evaluate-rank-filter in Table 1). Nevertheless, exactly evaluating tasks from either losses or gradients brings extra computational and annotation overhead. In LLMs, the hallucination stands as a lasting challenge, and the post-training phase sometimes relies on prohibitively intensive manual evaluation and precise human feedback such as millions of interactions, which serves for effective alignment 21, 22. In DR or Meta-RL, the agent has to interact with numerous MDPs to collect after-adaptation episodes and compute returns. All of these raise concerns about learning efficiency during optimization for robust adaptation, particularly when the machine learner employs a large backbone or interactions with environments are expensive.
Encouraged by the pressing real-world necessities, we delve into robust active task sampling, which can potentially cut off extra costs arising from task construction, intensive annotations, or computations when evaluating the machine learner’s adaptation to tasks. In the presence of any-shot adaptation, it is desirable to seek a task sampling strategy that requires fewer learning resources but retains more deployment benefits such as robust pattern recognition with foundation models and risk-averse sequential decision-making.
@@
Our brain can simulate the outcome of some unencountered scenarios based on continually cumulated experience without exactly putting actions into practice. The inherent cognitive system is also robust to handle extreme cases and energy efficient when developing diverse skills through an implicit task selection mechanism. Hence, it prompts us to learn to model the optimization process of any-shot adaptation and simultaneously adjust the task sampling strategy from the predicted outcome as feedback. A similar rationale is reflected in neuroscience 4, 23, 24; the prefrontal cortex resembles a meta-reinforcement learning system, implying the adaptation feedback also interleaves with the task query process. As a result, this work investigates the design of task robust sampling strategies with the guidance of a risk predictive model. We posit key insights in methodology: (i) Some variables, i.e., adaptation risk, are predictable during episodic learning, which is informative to rank task difficulties and steer task selection; (ii) Generative modeling task-specific adaptation risk enables capturing risk landscapes with quantified uncertainty, which aligns optimization processes with certain robustness principles.
To this end, we develop a new framework oriented to task episodic learning, namely Model Predictive Task Sampling (MPTS) to enable robust active task sampling in risk minimization. As displayed in Fig. 1a, our strategy exploits the historical risk information to construct a lightweight surrogate model referred to as the risk learner, foresighting adaptation risk across the task space for the task sampler and guiding the optimization of the adaptive machine learner. Specifically, we view actively selecting task batches as planning from past experience with the assistance of the risk learner in MPTS. Fig. 1b reveals the risk learner’s role in amortizing the expensive evaluation process on an arbitrary set of tasks conditioned on an adaptive machine learner, which further serves task difficulty scoring. The risk learner in Fig. 1c is in a variational autoencoder (VAE) structure 25 and engages in online adaptation risk generation via posterior inference. The acquisition function in Fig. 1d comprises worst-case performance and the predictive uncertainty into selection criteria to rate the task identifier candidate for filtering purposes.
MPTS also draws inspiration from active inference 26, which involves perception, action, and learning phases in a loop for uncertainty minimization about the planning environment. Similarly, this work conceptually characterizes task episodic learning by a risk generative model in Fig. 2b-c, and the subset selection from the task batch can be treated as online planning to derive a robust machine learner. Technically, MPTS specifies or infers the identifiers from the task distribution (see some instantiations in Fig. 2a) to establish mappings inside the risk learner and adopts the streaming variational inference (VI) in training the model. In addition, running acquisition functions in a larger identifier batch to pick up the task subset for the next iteration balances exploration (uncertainty minimization) and exploitation (worse-case robustness) in the task space. In primary, our proposed MPTS enjoys several benefits in robust optimization:
@@
    Adaptation Robustness. The optimization pipeline of MPTS can advance the machine learner’s adaptation robustness in the presence of severe task distribution shifts, such as subpopulation shifts and domain generalization;
@@
    Learning Efficiency. As the lightweight risk learner amortizes expensive evaluation parts, MPTS can diminish computational overhead, avoid unnecessary annotations, and encourage exploration in the task space;
@@
    Framework Versatility. Leveraging the optimization episodes, MPTS can work as a plug-play module to rank the task difficulties in optimization and allow seamless integration into robust zero-shot, few-shot, and many-shot learning.
This work performs extensive experiments in diverse task episodic learning scenarios, including few-shot regression, few-shot image classification with foundation models, Metra-RL, robotic DR, and prompt-tuning foundation models. Evidence indicates that incorporating MPTS into any-shot adaptation learning can significantly promote adaptation robustness in most scenarios. Compared with existing SOTA robust adaptation methods, MPTS raises no extra computation, memory usage, or interactions with environments and even accelerates the learning process under certain scenarios.
Figure 2: Fundamental Concepts: Task Identifiers, Episodic Learning and Probabilistic Graphical Models. a. The task distribution is uniform and defined over meaningful identifiers $\bm{\tau}$. For example, the amplitude and the phase $[a,b]$ specifies a sinusoid curve to complete with K-shot observed data points. Robots like Half-Cheetahs are trained to accomplish different locomotion tasks with varying masses and velocities. Some multimodal pattern recognition tasks’ identifiers are implicit but can be described from a reference model, e.g., text encoders in CLIP 1. b. The episodic learning in the task space resamples a batch of tasks from $p(\tau)$ to optimize in each iteration. c. Here, the generative model includes grey units as observed variables and white ones as unobservable. The solid directed lines describe the generative model 27. We use the dash-directed lines to indicate the recognition model and approximate inference within autoencoding variational Bayes 25.
@@ 2 Adaptation and Robustness
We denote a task sample by $\tau\sim p(\tau)$, with $\mathcal{T}$ the task space. The identifier, a real-value vector $\bm{\tau}$ illustrated in Fig. 2a, specifies the task $\tau$ in the distribution. Let the task-specific risk function be $\ell:\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}\times\Theta\mapsto\mathbb{R}$, which evaluates $\bm{\theta}$ parameterized machine learner’s adaptation performance on the task $\tau$. Take few-shot regression as an instance, $\mathcal{D}_{\tau}^{S}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=1}^{K}$ refers to the support dataset available for fast adaptation, e.g., finetuning to obtain the model $p_{\bm{\theta}}(\bm{y}|\mathcal{D}_{\tau}^{S},\bm{x})$; while $\mathcal{D}_{\tau}^{Q}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=K+1}^{K+N}$ denotes the query dataset to evaluate the after-adaptation machine learner’s performance, e.g., $\ell=-\frac{1}{N}\sum_{i=1}^{N}\ln p_{\bm{\theta}}(\bm{y}_{i}|\mathcal{D}_{\tau}^{S},\bm{x}_{i})$.
Given the condition $|\mathcal{D}_{\tau}^{S}|=\emptyset$, the risk function $\ell$ examines the zero-shot adaptation capability of a trained model. Otherwise, $\ell$ reflects the few-shot or many-shot adaptation risk (e.g., supervised finetuning over many examples). During optimization, we can write the history of the episodic task batch information as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$ with $\mathcal{B}$ the number of tasks in one batch, where $\bm{\theta}_{t}$ represents the machine learner’s parameter in the $t$-th iteration.
@@
The tuple set $\left\{\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$ involves the sampled task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$, the query and support dataset $\{\mathcal{D}_{\tau_{t,i}}:=\mathcal{D}_{\tau_{t,i}}^{S}\cup\mathcal{D}_{\tau_{t,i}}^{Q}\}_{i=1}^{\mathcal{B}}$, and the evaluated adaptation risk $\{\ell_{t,i}\}_{i=1}^{\mathcal{B}}$ in the $t$-th iteration. For convenience, we express the simplified risk history as $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$, which depends on $\bm{\theta}_{t}$.
2.1 Adaptation Risk Function and Task Filtering
@@
The task episodic learning setup optimizes the machine learning within $p(\tau)$. Our analysis is interested in the risk landscape in the task space as illustrated in Fig. 1d. Such a perspective emphasizes the interplay between the task identifier $\bm{\tau}$, the task-specific dataset $\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$ and the adaptation risk function $\ell$ conditioned on $\bm{\theta}$. To account for varying levels of adaptation, we delineate three distinct scenarios: zero-shot, few-shot, and many-shot learning. As finetuning is treated as a many-shot adaptation, we skip the instantiation below.
Zero-Shot Adaptation Risk Function.
@@
During training, we evaluate $\ell$ on the query dataset $\mathcal{D}_{\tau}^{Q}$ conditioned on the machine learner $\bm{\theta}$, i.e., $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$. With robotic DR 9 as an example, $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$ denotes the negative return of trajectories collected under the policy $\bm{\theta}$. This evaluation is commonly seen in zero-shot learning without support information.
Few-shot Adaptation Risk Function.
@@
The form of $\ell$ is specific to meta-learning methods. For instance, MAML 28 implements a bi-level optimization framework. In this context, $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})$ is written as $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta}_{\text{meta}}-\alpha\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{S};\bm{\theta}))$, where $\bm{\theta}_{\text{meta}}$ denotes the meta initialization, and the inside-bracket gradient term corresponds to finetuning the model tailored to $\tau$ with $\alpha$ the learning rate.
Task Filtering in Robust Adaptation Learning Pipelines.
Recent works 29, 19, 30 uncover compelling evidence that filtering some tasks to train under several criteria, e.g., ‘easy,’ ‘difficult,’ or ‘learnable’ in evaluation can enhance machine learners’ adaptation robustness and even improve task efficiency compared with uniform sampling strategies. The ‘difficult’ prioritization criteria aligns with minimizing conditional value-at-risk ($\text{CVaR}_{\alpha}$) 31, a metric in evaluating robustness 32, 16, 33. These methods operate under an Evaluate-Rank-Filter mode for the next model update in Table 1 and induce the batch filtering ratio $\hat{\alpha}=1-\frac{\mathcal{B}}{\hat{\mathcal{B}}}\in[0,1)$, which describes the fraction of discarded tasks in a sample batch.
Table 1: A Summary of Typical Risk Minimization Approaches. The previously mentioned ERM 34, GDRM 35, 36, 37, DRM 16, 33, 29, 30, and MPTS are included in comparison. In DRM, the filtering criteria is to discard proportional easy tasks in a batch.
        Principle         Robustness Concept         Optimization Pipeline
        ERM         No         Evaluate $\Rightarrow$ Update
        GDRM         Uncertainty Set         Evaluate $\Rightarrow$ Reweight $\Rightarrow$ Update
        DRM         Tail Task Risk         Evaluate $\Rightarrow$ Rank $\Rightarrow$ Filter $\Rightarrow$ Update
        MPTS (Ours)         Tail Task Risk with Uncertainty         Simulate $\Rightarrow$ Rank $\Rightarrow$ Filter $\Rightarrow$ Update
2.2 Risk Minimization Principles
Expected/Empirical Risk Minimization (ERM).
With the fixed $p(\tau)$, the principle follows the statistical learning theory 34 and minimizes the expectation of adaptation risk over the task space. As a result, we can have:
It draws batches with a random task sampler to optimize iteratively.
Group Distributionally Robust Risk Minimization (GDRM).
Such a principle 35 effectively improves robustness in distribution shifts and has shown positive effects on training foundation models 36, 37. It constructs a collection of uncertainty sets over tasks and results in the optimization objective as follows:
Distributionally Robust Risk Minimization (DRM).
We retain the notation of task robust works 16, 38, which terms the tail task risk minimization as DRM. It aims to improve the robustness of adaptation to the tail tasks over iteration. No explicit form exists as the tail task distribution is $\bm{\theta}$-dependent. The optimization objective is derived as the $\text{CVaR}_{\alpha}(\bm{\theta})$ 31:
where we write $p_{\alpha}(\tau;\bm{\theta})$ to express the $(1-\alpha)$ proportional worst case for easier formulation. Also note that when $\alpha$ approaches $1$, the problem degenerates to the worst-case risk minimization.

### v5

# Model Predictive Task Sampling for Efficient and Robust Adaptation

@@ Abstract
Foundation models have revolutionized general-purpose problem-solving, offering rapid task adaptation through pretraining, meta-training, and finetuning. Recent crucial advances in these paradigms reveal the importance of challenging task prioritized sampling to enhance adaptation robustness under distribution shifts. However, ranking task difficulties over iteration as a preliminary step typically requires exhaustive task evaluation, which is practically unaffordable in computation and data-annotation. This study provides a novel perspective to illuminate the possibility of leveraging the dual importance of adaptation robustness and learning efficiency, particularly in scenarios where task evaluation is risky or costly, such as iterative agent-environment interactions for robotic policy evaluation or computationally intensive inference steps for finetuning foundation models. Firstly, we introduce Model Predictive Task Sampling (MPTS), a framework that bridges the task space and adaptation risk landscape, providing a theoretical foundation for robust active task sampling. MPTS employs a generative model to characterize the episodic optimization process and predicts task-specific adaptation risk via posterior inference. The resulting risk learner amortizes the costly evaluation of task adaptation performance and provably approximates task difficulty rankings. MPTS seamlessly integrates into zero-shot, few-shot, and supervised finetuning settings. Empirically, we conduct extensive experiments in pattern recognition using foundation models and sequential decision-making. Our results demonstrate that MPTS significantly enhances adaptation robustness for tail or out-of-distribution (OOD) tasks and improves learning efficiency compared to state-of-the-art (SOTA) methods. The code is available at the project site https://github.com/thu-rllab/MPTS.
@@ 1 Introduction
Figure 1: Framework of MPTS in Adaptation Learning. a. The left is the standard optimization pipeline in Meta-Learning, DR, or SFT, with the machine learner a foundation model or an RL policy. The right is MPTS using a predict-then-optimize strategy, which integrates posterior inference, adaptation outcome simulation, and challenging task subset selection. [Snow: frozen models; Fire: updated models] b. With $\text{CVaR}_{\alpha}$ for DR or Meta-RL as an example, each iteration resamples $\hat{\mathcal{B}}$ MDPs for adaptive policies to interact and evaluate and then select $-\mathcal{B}$ worst MDP episodes for optimization. Dashed blue parts are what MPTS amortizes. c. The risk learner utilizes the risk history $H_{1:t}$ to train under a streaming VI framework. d. The risk learner simulates adaptation outcomes $p(\ell|\bm{\tau},H_{1:t};\bm{\theta}_{t})$ for $\hat{\mathcal{B}}$ candidate identifiers, computes acquisition scores, and selects the $-\mathcal{B}$ identifiers for the $(t+1)$-th iteration.
Generalization across diverse scenarios remains a central challenge in artificial general intelligence. The rise of generative AI offers a promising solution, driving the development of foundation models 1, 2, 3. Unlike traditional task-specific models, which might fail in new tasks, foundation models enable fast deployment across diverse scenarios without learning from scratch. Their rapid problem-solving stems from widely adopted adaptation learning paradigms, including pretraining, meta-learning, and supervised finetuning (SFT).
These paradigms train machine learners over a task distribution, consolidating past experience into problem-solving priors to handle unseen but related tasks in zero-shot or few-shot settings 1, 4. Each iteration samples a task batch, e.g., from a uniform distribution, and then executes the learning-to-adapt step (see Fig. 1/2b). Large language models (LLMs), for instance, treat episodic corpus datasets as tasks and perform in-context learning for adaptation 5. Similarly, in obtaining generalist robotic policies, decision-making environments, such as Markov decision processes (MDPs), are randomized for robots to perform policy optimization. These task distributions are typically determined by identifiers; e.g., in Fig. 2a, varying physics parameters configure different MDPs as tasks for domain randomization (DR) 6 and meta reinforcement learning (Meta-RL) 7.
@@
Distribution shifts 8, 9 are prevalent in real-world scenarios, making task adaptation robustness at test time increasingly critical 10, 11. In this context, task sampling strategies play a pivotal role, yet uniform sampling often causes catastrophic failures in risk-sensitive scenarios due to the undersampling of critical tasks. Two real-world applications highlight this case: (i) Tail tasks. In developing autonomous-driving systems, traffic accidents are rare in training datasets but disproportionately important for testing robustness 12. (ii) OOD tasks. Robots trained in controlled environments struggle in unstructured real-world settings, e.g., leading to errors in navigation and object manipulation. To improve robustness, challenging task prioritized sampling 12, 13, 14 has gained traction, where assessing task difficulty is central to robust optimization. These methods 15, 16, 14, 13, 17, 18 evaluate, rank, and prioritize difficult tasks for iterative optimization (see Fig. 1b). However, precisely evaluating tasks—via losses, human annotations, or gradients—incurs high computational costs. For instance, in LLM alignment, task evaluation through SFT requires extensive forward passes, while preference optimization consumes millions of expert annotations 19. Similarly, in DR and Meta-RL, agents must interact with numerous MDPs to collect post-adaptation episodes and compute returns. These challenges uncover the urgent need for more efficient learning strategies when enhancing robust adaptation, particularly when deploying foundation models or when environment interactions are costly.
Motivated by the above pressing demands, we dive into robust active task sampling, a paradigm that has the potential to eliminate unnecessary costs associated with task construction, intensive annotations, or computational overhead during the evaluation of a machine learner’s adaptation to tasks. In scenarios involving zero-shot learning, few-shot learning, or SFT, we aim to develop a task sampling strategy that requires fewer learning resources but retains more deployment benefits such as adaptation robustness in pattern recognition with foundation models and risk-averse sequential decision-making.
@@
Note that our brain is energy-efficient and simulates the outcome of decision-making in unencountered scenarios from accumulated experience, without actual trials. This capability arises from mechanisms like implicit information gating and active task selection 20, 21, 22. Inspired by this, we propose a model-based optimization approach for adaptive learning, dynamically adjusting task sampling strategies using predicted outcomes as feedback. This work explores the design of risk predictive models, referred to as risk learners, for robust task sampling based on two key insights: (i) Adaptation risk is probably predictable in episodic learning, providing a basis for task difficulty ranking and selection; (ii) Generative modeling of adaptation risk captures risk landscapes with quantified uncertainty, aligning optimization with robustness principles.
To this end, we introduce Model Predictive Task Sampling (MPTS), a framework for risk-aware task selection. As shown in Fig. 1a, MPTS leverages historical risk to train a lightweight risk learner, which forecasts adaptation risks across the task space to guide the task sampler and optimize the adaptive machine learner. This way amortizes expensive tasks’ evaluation for ranking their difficulty to select subset (see Fig. 1b). The risk learner in Fig. 1c adopts a variational autoencoder (VAE) 23 structure, generating adaptation risk estimates via posterior inference 24. Finally, the acquisition function in Fig. 1d integrates worst-case performance and predictive uncertainty into the rule of subset selection.
MPTS also draws inspiration from active inference 25, which operates through a loop of perception, action, and learning to minimize uncertainty about the planning environment. Here, subset selection from the task batch can be viewed as online planning to derive a robust machine learner. Technically, MPTS specifies or infers identifiers from the task distribution (see examples in Fig. 2a) to establish mappings between identifiers and adaptation risk. It employs streaming variational inference (VI) for risk learner training. Furthermore, by simulating adaptation outcomes in a larger identifier batch for subset selection, MPTS balances exploration (uncertainty minimization) and exploitation (worst-case robustness) across the task space. In primary, our proposed MPTS enjoys several benefits in practice:
@@
    Adaptation Robustness. The optimization pipeline of MPTS can advance the machine learner’s adaptation robustness under severe task distribution shifts, such as tail or OOD task scenarios;
@@
    Learning Efficiency. Constructing the lightweight risk learner to amortize expensive evaluation, MPTS can diminish computational overhead, avoid unnecessary annotations, and promote efficient exploration in the task space;
@@
    Framework Versatility. Learning from risk histories, MPTS serves as a plug-play module to rank the task difficulties in optimization and allows seamlessly integration into robust zero-shot or few-shot learning and SFT.
This work evaluates MPTS across few-shot regression, image classification with foundation models, Meta-RL, robotic DR, and prompt-tuning foundation models. Empirical results demonstrate MPTS’s outstanding adaptation robustness across diverse scenarios. Compared to SOTA robust adaptation methods, MPTS significantly reduces computational overhead, memory usage, and environment interactions while, in some cases, accelerating learning.
Figure 2: Fundamental Concepts: Task Identifiers, Episodic Learning and Probabilistic Graphical Models. a. The task distribution is uniform and defined over meaningful identifiers $\bm{\tau}$. For example, the amplitude and the phase $[a,b]$ specifies a sinusoid curve to complete with K-shot observed data points. Robots like Half-Cheetahs are trained to accomplish different locomotion tasks with varying masses and velocities. Some multimodal pattern recognition tasks’ identifiers are implicit but can be described from a reference model, e.g., text encoders in CLIP 1. b. The tail task generalization corresponds to $\text{CVaR}_{\alpha}$, i.e., the integral of tail task risk values in red. In OOD generalization, this work prompt-tunes CLIP on ImageNet 26 to test on ImageNet-S 27. c. Here, the generative model includes grey units as observed variables and white ones as unobservable. The solid directed lines describe the generative model 28. We use the dash-directed lines to indicate the recognition model and approximate inference within autoencoding variational Bayes 23.
@@ 2 Adaptation and Robustness
We represent a task sample by $\tau\sim p(\tau)$, with $\mathcal{T}$ denoting the task space. Each task $\tau$ within the distribution is specified by an identifier, a real-valued vector $\bm{\tau}$, as illustrated in Fig. 2a. The task-specific risk function $\ell:\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}\times\Theta\mapsto\mathbb{R}$ evaluates the adaptation performance of a machine learner $\bm{\theta}$ on $\tau$. For example, in regression, the support dataset $\mathcal{D}_{\tau}^{S}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=1}^{K}$ enables rapid adaptation to obtain the model $p_{\bm{\theta}}(\bm{y}|\mathcal{D}_{\tau}^{S},\bm{x})$; while the query dataset $\mathcal{D}_{\tau}^{Q}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=K+1}^{K+N}$ is used for post-adaptation evaluation as risk $\ell=-\frac{1}{N}\sum_{i=1}^{N}\ln p_{\bm{\theta}}(\bm{y}_{i}|\mathcal{D}_{\tau}^{S},\bm{x}_{i})$.
@@
If $|\mathcal{D}_{\tau}^{S}|=\emptyset$, $\ell$ measures zero-shot adaptation; otherwise, it reflects few-shot adaptation risk. In SFT, each sample $(\bm{x},\bm{y})\in\mathcal{D}_{\text{SFT}}$ is treated as a task. The episodic task batch history is defined as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$, where $\mathcal{B}$ is the task batch size and $\bm{\theta}_{t}$ represents the machine learner’s parameter in $t$-th iteration. The tuple set $\left\{\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$ includes the sampled task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$, the support and query dataset $\{\mathcal{D}_{\tau_{t,i}}:=\mathcal{D}_{\tau_{t,i}}^{S}\cup\mathcal{D}_{\tau_{t,i}}^{Q}\}_{i=1}^{\mathcal{B}}$, and the evaluated adaptation risk $\{\ell_{t,i}\}_{i=1}^{\mathcal{B}}$. For simplicity, the risk history is expressed as $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$, which depends on $\bm{\theta}_{t}$.
Adaptation Risk Function.
@@
The learning setup optimizes the machine learner within $p(\tau)$. Our analysis is interested in the risk landscape in the task space as illustrated in Fig. 1d. Such a perspective emphasizes the interplay between the task identifier $\bm{\tau}$, the task-specific dataset $\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$ and the adaptation risk function $\ell$ conditioned on $\bm{\theta}$. Regarding adaptation performance, we mainly examine zero-shot learning, few-shot learning, and SFT scenarios.
@@
Zero-Shot Adaptation. During training, we evaluate $\ell$ on the query dataset $\mathcal{D}_{\tau}^{Q}$ conditioned on the machine learner $\bm{\theta}$, i.e., $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$. With robotic DR 6 as an example, $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$ denotes the negative return of trajectories collected under the policy $\bm{\theta}$ in MDP $\tau$. This setup is without support information.
@@
Few-Shot Adaptation. The form of $\ell$ is specific to meta-learning methods. For instance, MAML 29 implements a bi-level optimization framework. In this case, $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})$ is written as $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta}_{\text{meta}}-\alpha\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{S};\bm{\theta}))$, where $\bm{\theta}_{\text{meta}}$ denotes the meta initialization, and the inside-bracket term corresponds to finetuning $\bm{\theta}_{\text{meta}}$ tailored to $\tau$ with $\alpha$ the learning rate.